OpenStudy (anonymous):

In a coordinate system points A(-2,3) and B(6,-1) have a midpoint of x.What is the midpoint between A and x? A.(0,0) B.(2,1) C.(0,2) D.(4,2) E.(-1,1.5)

OpenStudy (anonymous):

well first you find the midpoint x, which is \[\frac{ B-A }{ 2 }\]

OpenStudy (anonymous):

so the x coordinate is \[\frac{ 6-(-2) }{ 2 }\]

OpenStudy (anonymous):

and y is \[\frac{ -1-3 }{ 2 }\]

OpenStudy (anonymous):

aaa my bad, it should be plus not minus

OpenStudy (anonymous):

so x is \[\frac{ -2+6 }{ 2 }\] and y is \[\frac{ -1+3 }{ 2 }\]

OpenStudy (anonymous):

so x = (2,1)

OpenStudy (anonymous):

And now using the same formula you find the midpoint between x and A

OpenStudy (anonymous):

I get C. (0,2)

OpenStudy (anonymous):

Thanks for the help :)

OpenStudy (anonymous):

No probs :D

OpenStudy (anonymous):

Helps alot when i have someone else helping me i couldnt understand it i was like omgg how do i do this XD

OpenStudy (anonymous):

hahha yeah, I get that sometimes too :p That's why this is a cool place